# Unifying KV Cache Compression for Large Language Models with LeanKV

### v1

@@ 3.2. Dynamic Sparsity Patterns
Figure 2. Number of critical tokens per layer in Llama3-8B to preserve 95% of the total attention score.
@@
Figure 3. Number of critical tokens per KV head in Llama3-8B to preserve 95% of the total attention score.
3.3. Unifying Quantization and Pruning
Quantization compresses the KV cache by representing feature values with a reduced bit width, whereas pruning can be viewed as an extreme form of quantization: effectively storing feature values with zero bits. By using attention scores to measure the significance of tokens, we can unify the spectrum of KV cache compression by storing progressively less critical tokens at decreasing precision levels, ultimately leading to the pruning of the least critical tokens.
To investigate the feasibility of this unified approach, we experiment with a simple design: quantizing both the key and value vectors of significant tokens to 8 bits while pruning insignificant ones. Figure 4 compares the performance of the FP16 baseline, 8-bit quantization, pruning, and the combination of 8-bit quantization with pruning on the CNN/Daily Mail summarization dataset using the Llama3-8B model. For both pruning and the combined approach, approximately 20% of insignificant tokens are pruned. The ROUGE-1 score, which measures the fidelity of the LLM-generated responses to reference answers, is nearly indistinguishable among the four approaches. However, the combination of 8-bit quantization and pruning achieves the lowest memory usage. The improved efficiency of this combined approach motivates us to further explore the spectrum of KV cache compression by introducing varying levels of precision to accommodate tokens of differing significance.
Figure 4. Rogue-1 score and normalized KV cache memory usage of the Llama3-8B model on CNN/DailyMail.
3.4. Key Findings and Implications
In this section, we outline the key findings that motivate the design of LeanKV. Specifically, we demonstrate that within individual tokens, key vectors have a greater impact than value vectors, inspiring us to propose storing key vectors at higher precision than value vectors. Beyond that, the significance of different tokens, as indicated by softmax scores, exhibits dynamic sparsity patterns across both heads and requests. These patterns necessitate flexible memory management on a per-head and per-request basis. Lastly, we show that unifying quantization and pruning can outperform either approach when used in isolation.
These findings motivate us to further explore the spectrum of KV cache compression, on a per-head and per-request basis to achieve the maximum memory efficiency. Specifically, each attention head should be allowed to store tokens at varying precision levels and to prune the least important ones, according to the sparsity patterns of individual requests. Additionally, within each token, the key and value vectors should be stored in different precisions. However, this unified approach presents significant challenges on memory management systems due to the resulting irregular memory layout, where the KV cache of each head and request consists of varying number of tokens stored at different precisions.
4. Unified Compression Algorithm
@@
In this section, we present the design of LeanKV’s unified compression algorithm, which leverages the insights from Section 3 to optimize memory efficiency. To account for the greater impact of key vectors compared to value vectors, we propose Hetero-KV, a strategy that quantizes keys at a higher precision than values. For instance, key and value vectors can be quantized to 8 and 4 bits (K8V4), or to 4 and 2 bits (K4V2). Additionally, we propose per-head memory management tailored to dynamic sparsity patterns. Rather than imposing a fixed memory budget, each head adaptively determines its memory allocation by quantizing tokens at various precision levels or pruning, based on the attention score distributions of individual requests. Specifically, tokens are categorized into three significance classes: the most important tokens are quantized at high precision (e.g., K8V4), moderately important tokens at low precision (e.g., K4V2), and the least important tokens are pruned.
@@
Next, we describe the unified compression algorithm in greater detail, addressing the prompt phase and the generation phase separately. Notably, in both phases, compression is conducted on a per-head and per-request basis.
Prompt Phase. In the prompt phase, key and value vectors for all tokens in the prompt are computed. The compression algorithm then determines the appropriate precision for storing each token based on its significance. The significance of the $i^{th}$ token is calculated by averaging the $N-i$ attention scores it receives from subsequent tokens, where $N$ denotes the sequence length. This average is then normalized to ensure that the total significance across all tokens sums to one.
To prevent premature compression, the algorithm reserves a recent window of size $W$, where all tokens are quantized at a high precision. $W$ is predefined and typically set to 64. The remaining tokens are then sorted by significance. The least important $N_{p}$ tokens, whose cumulative significance is below the threshold $T_{l}$, are pruned. The next $N_{l}$ tokens, with a cumulative significance less than $T_{h}$, are quantized to low precision, while the most important $N_{h}$ tokens are quantized to high precision. Here, $N_{p}+N_{l}+N_{h}+W=N$, and $T_{l}$ and $T_{h}$ are predefined thresholds determined by offline profiling. As a result of this mixed-precision compression, the KV cache is conceptually divided into two parts: a high precision section $KV_{h}$ and a low precision section $KV_{l}$.
@@
The detailed procedure is outlined in Algorithm 1 and is governed by two predefined thresholds: the high precision threshold $T_{h}$ and the low precision threshold $T_{l}$. If the significance of $t_{c}$ is at least $T_{h}$, it is quantized at high precision and added to $KV_{h}$, the high precision section of the KV cache (line 6). Subsequently, the least significant token in $KV_{h}$ is identified as the victim token $t_{v}$ for more aggressive compression (line 7). If $t_{v}$’s significance exceeds $T_{h}$, it remains in $KV_{h}$; if it falls between $T_{h}$ and $T_{l}$, $t_{v}$ is re-quantized to low precision and moved to $KV_{l}$, the low precision section of the KV cache (line 9); otherwise, $t_{v}$ is pruned. Similarly, if $t_{c}$’s significance falls between $T_{h}$ and $T_{l}$, it is quantized to low precision and added to $KV_{l}$ (line 14). The least important token in $KV_{l}$ is designated as the victim $t_{v}$ (line 15), and is further pruned if its significance falls below $T_{l}$ (line 17).
Notably, the proposed unified compression algorithm is extensible beyond three precision levels, and we adopt three precision levels to minimize metadata overhead and improve system efficiency, as detailed in the following section.
1:  Input: Thresholds $T_{h}$ $T_{l}$; Candidate token $t_{c}$
2:  Input: High & low precision $P_{h}$ $P_{l}$
@@
5:  if Score($t_{c}$) $\geq$ $T_{h}$ then
@@
10:     else if Score($t_{v}$) ¡ $T_{l}$ then
@@
13:  else if Score($t_{c}$) $\geq$ $T_{l}$ then
@@
Algorithm 1 Unified compression (generation phase)
5. System Design
In this section, we present the design of the LeanKV system that leverages the proposed unified KV cache compression algorithm to boost LLM inference efficiency. We first discuss the challenges that the compression algorithm poses to system performance, and then present novel techniques to tackle these challenges.
5.1. Challenges
@@
The Need for Flexible Paging. In PagedAttention (kwon2023efficient, ), all tokens are stored at the same precision, making a fixed page format sufficient. However, Hetero-KV and mixed-precision quantization introduce varying precision levels across different tokens, rendering a fixed page format inadequate. Specifically, a fixed page format requires conservatively allocating high-precision slots for all tokens regardless of their actual precisions, leading to considerable memory wastage. Suppose memory is allocated to accommodate the highest precision K8V4, a token with K4V2 precision would waste 50% of the memory. Additionally, such a fixed page format results in misaligned memory accesses, which hinder memory bandwidth utilization and reduce overall computational efficiency. To fully realize the benefits of Hetero-KV and mixed-precision quantization, we need a more flexible paging design that can dynamically adapt to varying precision levels, while ensuring compact memory usage and efficient memory accesses.
@@
Increased Memory Management Overhead. The per-head KV cache compression algorithm in LeanKV introduces both significant latency and metadata overhead in memory management. In PagedAttention, the memory management complexity is $O(\#requests)$ regardless of the number of heads, as it partitions memory uniformally across heads and assigns identical page IDs to each head. In contrast, LeanKV must handle varying numbers of high-precision and low-precision tokens across different heads, which increases the complexity to $O(\#requests\times\#heads)$. If not carefully optimized, this overhead could overshadow the performance gains from cache compression. Moreover, to track both high-precision and low-precision tokens, a straightforward approach would involve using separate data structures for each precision, resulting in increased metadata overhead.
5.2. On-GPU Memory Management
@@
A detailed analysis of the memory management process reveals opportunities for parallelization and acceleration. Specifically, memory management for the unified compression algorithm consists of two phases: planning and coordination. In the planning phase, each head independently determines its memory allocation requirements. The subsequent coordination phase synchronizes these per-head requirements and maps them to the GPU’s physical memory. For instance, assume two attention heads require 16 and 32 bytes, respectively, as determined in the planning phase, with all GPU physical memory available. Then in the coordination phase, physical memory bytes 0 to 15 are allocated to the first head, and bytes 16 to 48 to the second. The primary source of increased memory management complexity lies in the planning phase, which is perfectly parallel and well-suited to the GPU’s parallel compute capabilities. Parallelizing the coordination phase is more challenging, as it requires synchronization across heads. However, we observe that this phase can be parallelized effectively via parallel prefix sum (nakano2012prefix-sum, ).
@@
Thus, we design an on-GPU memory management system to efficiently handle per-head dynamic memory allocation and recycling. This system introduces three GPU-resident data structures that collectively tackle the aforementioned challenges: the $(1)$ unified paging to enable flexible paging, $(2)$ the circular free page list to efficiently execute the coordination phase in memory management, and $(3)$ the bidirectional page table to minimize metadata overhead for tracking high-precision and low-precision tokens.
@@
Unified Paging. We partition the available GPU memory into evenly sized pages, each configured upon allocation to store tokens at a given precision. Each unified page consists of six segments: quantized keys, quantization metadata for keys, quantized values, quantization metadata for values, token scores, and positions. The quantization metadata includes scales and zero points for the key and value vectors. The number of tokens stored per page is adjusted according to the quantization configuration, ensuring compact memory usage. Unified paging allows flexible handling of tokens with varying precisions while maintaining a consistent memory layout, effectively supporting our unified compression algorithm. Furthermore, by consolidating keys, values, and their metadata into a single structure, unified paging enhances data locality and eliminates the need for scattered lookups, improving memory access efficiency during attention computation.
@@
Circular Free Page List. To facilitate efficient page allocation and recycling, we manage all available page IDs in a centralized, GPU-resident data structure called circular free page list. The list tracks free pages using a pair of start and end pointers, which wrap to the beginning of the list upon reaching the end, maintaining a circular structure. When a page is allocated, the start pointer increments to the next available page ID; when a page is freed, the end pointer increments to add the freed page ID back to the list.
@@
Notably, both the available and unavailable regions in the list remain contiguous, enabling the coordination phase in memory management to be parallelized via parallel prefix sum. Specifically, after each head determines the number of pages to be allocated or freed in the planning phase, a parallel prefix sum operation computes a unique offset for each head relative to the start or end pointer, ensuring non-conflicting operations within the list. For memory allocation, each head concurrently retrieves its new page IDs from its designated region in the list, with the start pointer incremented by the cumulative number of pages required across all heads. Similarly, for memory recycling, each head concurrently writes freed page IDs to it designated region, with the end pointer incremented by the total number of pages released.
@@
Bidirectional Page Table. Similar to PagedAttention, LeanKV maintains a page table for each attention head, mapping requests to physical memory. Each entry in the page table is a list of page IDs allocated to the head for a specific request. To avoid the doubled metadata overhead associated with maintaining separate page tables for high- and low-precision pages, LeanKV introduces a unified, GPU-resident data structure called the bidirectional page table. This structure efficiently supports the use of three precision levels for KV cache compression, as proposed in Section 4, minimizing the metadata overhead. In each entry of the bidirectional page table, high-precision page IDs grow from the left side of the list, while low-precision page IDs grow from the right, dynamically adapting to the precision requirements of the workload. The length of each page table entry is determined by the maximum sequence length divided by the tokens per high-precision page, ensuring no overflow since low-precision pages always contain more tokens than high-precision ones. This unified approach not only minimizes metadata overhead but also eliminates the need for separate lookups based on precision levels, thereby enhancing memory access efficiency during attention computations. The memory overhead of the bidirectional page table is minimal: for example, with a batch size of 128 on Llama-2-7B, which has 32 layers and 32 attention heads per layer, the total size of all bidirectional page tables is only 64 MB. By contrast, the KV cache for a single request occupies 2 GB.
5.3. Memory Management Workflow
@@
In this section, we delve into LeanKV’s memory management workflow, demonstrating how the proposed data structures interact to efficiently support unified KV cache compression. Notably, this memory management workflow is executed once per inference step for all requests in the batch and across all layers in the LLM. This design choice ensures sufficient parallelism for efficient GPU execution and amortizes the associated GPU kernel launch overheads.
Figure 5. Memory management flow in the prompt phase.
@@
Prompt Phase. Figure 5 illustrates the memory management workflow during the prompt phase using an example request with eight tokens. In this setting, one high-precision page can store two tokens, and one low-precision page can store four. Since the exact number of high- and low-precision pages required by each head is unknown prior to running the compression algorithm, we conservatively allocate four unified pages per head, assuming that all tokens will be stored at high precision. Each head then independently applies the KV cache compression algorithm to determine which tokens to store at high-precision, low-precision, or to prune. This step constitutes the planning phase, during which each head calculates its specific memory needs based on its unique sparsity pattern. In this example, the first head uses one high-precision page, while the second head uses two; both heads also require one low-precision page. High-precision pages are allocated from the left side of the bidirectional page table, while low-precision pages are allocated from the right side. Following the planning phase, any unused pages are marked for recycling, triggering the coordination phase. Here, we employ a parallel prefix-sum-based approach, as described in the previous section, to facilitate efficient and conflict-free page recycling across heads.
@@
6. Implementation
We implement LeanKV on top of vLLM, comprising 4.5K lines of CUDA/C++ code and 9K lines of Python. We first outline the architecture of LeanKV and then detail our custom GPU attention kernel, designed to efficiently support the unified cache compression algorithm.
6.1. LeanKV Architecture
The architecture of LeanKV is illustrated in Figure 6. At each inference step, the scheduler batches as many requests as possible within the available GPU memory to maximize throughput, sending the selected requests to all workers. Each GPU hosts one worker, responsible for executing a segment of the model. LeanKV utilizes tensor parallelism (shoeybi2019megatron, ), where each worker handles a partition of each matrix multiplication and a subset of the attention heads. Each worker includes a dedicated memory manager to oversee the KV cache for its assigned attention heads, using the method described in Section 5. Due to dynamic sparsity patterns, memory requirements vary per head for each request, leading to an irregular memory layout. Additionally, each worker incorporates an execution engine for model computation. To support unified KV cache compression, the execution engine integrates a KV compressor and a custom GPU attention kernel. After computing key and value vectors, the KV compressor is invoked to compress them using the algorithm in Section 4, storing the results in the KV cache. The custom GPU attention kernel then efficiently computes the attention output using the compressed KV cache.
Figure 6. LeanKV architecture.
6.2. Efficient Attention Kernel
We develop a custom GPU attention kernel that efficiently supports per-head dynamic sparsity and Hetero-KV quantization, fully harnessing the reduction in memory access volume and translating it into accelerated performance.
Overall, we assign each CUDA thread block to process a single attention head per sequence. However, mixed-precision quantization introduces potential load imbalance, as high-precision tokens demand more bandwidth, which is the bottleneck of attention, than low-precision ones. Thanks to the unified paging mechanism that stores tokens of the same precision together in a page, we mitigate this issue by having thread warps iterate over high-precision pages first, followed by low-precision pages, ensuring that threads within a warp remain load-balanced. Each page is processed in two phases: dot product between the query and keys to derive attention scores, and weighted sum of values. These two phases follow different computation patterns: reduction over the feature dimension and token dimension, respectively. To ensure coalesced and vectorized memory accesses for both phases, we design tailored data layouts and parallelization strategies, elaborated in the following paragraphs.

### v2

@@ 3.3 Dynamic Sparsity Patterns
Figure 3: Number of critical tokens per layer in Llama3-8B to preserve 95% of the total attention score.
@@
Figure 4: Number of critical tokens per KV head in Llama3-8B to preserve 95% of the total attention score.
3.4 Key Findings and Implications
Our findings reveal three critical levels of differentiation within the KV cache:
  • Keys exert a broader impact on attention computation than values, motivating differentiated precisions for keys and values.
  • Tokens vary in importance as reflected by their attention scores, motivating a fine-grained, hierarchical compression strategy.
  • Attention sparsity patterns vary across requests and attention heads, requiring dynamically managing memory resources on a per-request and per-head basis.
Collectively, these insights guide the design of LeanKV, which integrates differentiated precision assignments for keys and values, hierarchical token importance handling, and per-request and per-head memory management.
4 KV Compression Policy
@@
We present the KV compression policy of LeanKV that addresses the three levels of differentiation within the KV cache. First, to reflect the greater impact of keys on attention computation than values, we propose quantizing keys at a higher precision than values. For example, keys and values can be quantized to 8 and 4 bits (K8V4), or 4 and 2 bits (K4V2). Second, to account for the varying importance of tokens, we introduce a hierarchical compression strategy that classifies tokens into three significance levels: the most important tokens are quantized at high precision (e.g., K8V4), moderately important tokens at lower precision (e.g., K4V2), and the least important tokens are pruned. Finally, to address the dynamic sparsity patterns across requests and attention heads, we propose an adaptive memory management approach. Rather than imposing a fixed memory budget, our approach allows each attention head to determine its memory requirements dynamically based on its specific sparsity pattern.
@@
Next, we describe the KV compression policy in greater detail, addressing the prompt phase and the generation phase separately. In both phases, compression is applied per-request and per-head, ensuring that memory usage is tailored to the specific sparsity pattern of each request and attention head.
Prompt Phase. In the prompt phase, key and value vectors for all tokens in the prompt are computed. The compression policy then determines the appropriate precision for storing each token based on its significance. The significance of the $i^{th}$ token is calculated by averaging the $N-i$ attention scores it receives from subsequent tokens, where $N$ denotes the sequence length. In the case of GQA and MHA, scores from all attention heads associated with the KV head are aggregated using the maximum operation. To mitigate premature compression, the most recent $W$ tokens are always quantized at high precision, where $W$ is typically set to 64. For the remaining tokens, the policy determines the precision level of the $i^{th}$ token by comparing its significance score to the theoretical average $\frac{1}{i}$, based on the intuition that tokens with scores below average are less critical. Specifically, the $i^{th}$ token is quantized at high precision if its significance score exceeds $\frac{\alpha_{h}}{i}$, at low precision if its score lies within the interval $[\frac{\alpha_{l}}{i},\frac{\alpha_{h}}{i}]$, and is pruned otherwise. The parameters $\alpha_{l}$ and $\alpha_{h}$, which define the thresholds for high- and low-precision quantization, are determined offline through profiling on a calibration dataset. As a result of this hierarchical compression, the KV cache is conceptually divided into two parts: a high-precision section $KV_{h}$ and a low-precision section $KV_{l}$.
@@
The detailed procedure is outlined in Algorithm 1, following the same intuition and parameters as the prompt phase. Given the sequence length $N$, $t_{c}$ is quantized at high precision and added to $KV_{h}$, the high precision KV cache (line 6), if its significance exceeds $\frac{\alpha_{h}}{N}$. Subsequently, the least significant token in $KV_{h}$ is identified as the victim token $t_{v}$ for more aggressive compression (line 7). If $t_{v}$’s significance exceeds $\frac{\alpha_{h}}{N}$, it remains in $KV_{h}$; if it falls within $[\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, $t_{v}$ is re-quantized to low precision and moved to $KV_{l}$, the low precision KV cache (line 9); otherwise, $t_{v}$ is pruned. Similarly, if $t_{c}$’s significance lies within $[\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, it is quantized to low precision and added to $KV_{l}$ (line 14). The least important token in $KV_{l}$ is designated as the victim $t_{v}$ (line 15), and is further pruned if its significance falls below $\frac{\alpha_{l}}{N}$ (line 17).
Discussion. The proposed KV compression policy is highly extensible, allowing for the potential use of more than three precision levels. We adopt three precision levels primarily to minimize metadata overhead and improve system efficiency, as will be detailed in the following section. Additionally, we use a shared set of thresholds for all attention heads, as our empirical studies demonstrate that this approach is sufficient to capture the varying sparsity patterns across different heads. Notably, LeanKV can efficiently support more flexible compression policies, such as tuning thresholds for each head individually, which could potentially further optimize the balance between model accuracy and memory efficiency. We leave the exploration of such flexible policies for future work.
1:  Input: Parameters $\alpha_{h}$, $\alpha_{l}$; High & low precision $P_{h}$ $P_{l}$
2:  Input: Candidate token $t_{c}$; Sequence length $N$
@@
5:  if Score($t_{c}$) $\geq$ $\frac{\alpha_{h}}{N}$ then
@@
10:     else if Score($t_{v}$) < $T_{l}$ then
@@
13:  else if Score($t_{c}$) $\geq$ $\frac{\alpha_{l}}{N}$ then
@@
Algorithm 1 KV compression policy (generation)
5 Memory Management
In this section, we describe LeanKV’s memory management mechanism which efficiently supports the three levels of differentiation in the KV cache. We first highlight the challenges posed by these differentiations in memory management, and then present parallel KV compaction, a novel memory management technique that effectively tackles the challenges.
5.1 Challenges
@@
The Need for Flexible Paging. In PagedAttention [32], all tokens are stored at the same precision, allowing for a fixed page format. However, the differentiation of key and value vectors, along with varying token importance, introduces multiple precision levels both within individual tokens and across different tokens, making a fixed page format insufficient. Specifically, a fixed page format requires conservatively allocating high-precision slots for all tokens regardless of their actual precisions, leading to considerable memory wastage. Suppose memory is allocated to accommodate the highest precision K8V4, a token with K4V2 precision would waste 50% of the memory. Worse, such a fixed page format results in misaligned memory accesses, which hinder memory bandwidth utilization and reduce overall computational efficiency.
@@
Substantial KV Compaction Overhead. In PagedAttention, memory management complexity is $O(\#requests)$, as memory is partitioned uniformly across attention heads, with identical page IDs assigned to each head. In contrast, LeanKV must handle varying numbers of high- and low-precision tokens across different heads, mapping these irregular memory requirements to physical memory. We term this process KV compaction, whose complexity is $O(\#requests\times\#heads)$, significantly higher than that of PagedAttention. If KV compaction were performed sequentially on CPUs, its overhead could overshadow the performance gains from KV cache compression. Moreover, to track both high-precision and low-precision tokens, a straightforward approach would involve using separate data structures for each precision, resulting in increased metadata overhead.
5.2 Parallel KV Compaction
@@
A detailed analysis of the KV compaction process reveals opportunities for parallelization and performance optimization. KV compaction can be divided into two phases: planning and coordination. In the planning phase, each head independently determines its memory allocation requirements. The subsequent coordination phase synchronizes these per-head requirements and maps them to the GPU’s physical memory. For instance, consider two attention heads require 16 and 32 bytes, respectively, as determined in the planning phase, with all GPU physical memory available. Then in the coordination phase, physical memory bytes 0 to 15 are allocated to the first head, and bytes 16 to 48 to the second. The primary source of increased memory management complexity lies in the planning phase, which is perfectly parallel and well-suited to the GPU’s parallel compute capabilities. Parallelizing the coordination phase is more challenging, as it requires synchronization across heads. However, we observe that this phase can be parallelized effectively via parallel prefix sum [44], provided that the free memory region is contiguous.
@@
Thus, we propose parallel KV compaction, a novel management technique that efficiently performs per-head dynamic memory allocation and recycling in parallel directly on the GPU. Parallel KV compaction is enabled by three GPU-resident data structures collectively: the $(1)$ unified pages to enable flexible paging, $(2)$ the circular free page list to efficiently parallelize memory management, and $(3)$ the bidirectional page table to minimize metadata overhead for tracking tokens of differentiated precisions.
@@
Unified Pages. Unified pages abstract away the complexity of differentiated precisions both within individual tokens and across different tokens, thereby simplifying the implementation of parallel KV compaction. Specifically, we partition the available GPU memory into evenly sized pages, each configured upon allocation to store tokens at a given precision. Each unified page is organized into six segments: quantized keys, quantization metadata for keys, quantized values, quantization metadata for values, token scores, and positions. The quantization metadata includes scales and zero points for the key and value vectors. The number of tokens stored per page is adjusted according to the quantization configuration, ensuring compact memory usage. Furthermore, by consolidating keys, values, and their metadata into a single structure, unified pages enhance data locality and eliminate the need for scattered lookups, improving memory access efficiency during attention computation.
Circular Free Page List. The circular free page list serves as the cornerstone of parallel KV compaction, facilitating the parallelization of memory allocation and recycling by maintaining both free and used pages in contiguous regions.
@@
This centralized, GPU-resident data structure contains all Page IDs and tracks free pages through a pair of pointers: a start pointer for allocations and an end pointer for recycling. These pointers wrap around to the beginning of the list upon reaching the end, forming a circular structure. When a page is allocated, the start pointer advances to the next available Page ID; when a page is freed, the end pointer advances to add the released Page ID back into the list. Both the available and unavailable regions in the list remain contiguous, enabling the coordination phase in memory management to be parallelized via parallel prefix sum.
@@
In parallel KV compaction, after each head determines the number of pages to be allocated or freed, a parallel prefix sum operation computes a unique offset for each head relative to the start or end pointer, ensuring non-conflicting operations within the list. For memory allocation, each head concurrently retrieves its new page IDs from its designated region in the list, with the start pointer incremented by the cumulative number of pages required across all heads. Similarly, for memory recycling, each head concurrently writes freed page IDs to its designated region, with the end pointer incremented by the total number of pages released.
@@
Bidirectional Page Table. Similar to PagedAttention, LeanKV maintains a page table for each attention head, mapping each request to its corresponding list of page IDs. To avoid the doubled metadata overhead associated with maintaining separate page tables for high- and low-precision pages, LeanKV introduces a unified, GPU-resident data structure called the bidirectional page table. This structure efficiently supports the use of three precision levels for KV cache, as proposed in Section 4, minimizing the metadata overhead. In each entry of the bidirectional page table, high-precision page IDs grow from the left side of the list, while low-precision page IDs grow from the right, dynamically adapting to the precision requirements of the workload. The length of each page table entry is determined by the maximum sequence length divided by the tokens per high-precision page, ensuring no overflow since low-precision pages always contain more tokens than high-precision ones. This unified approach not only minimizes metadata overhead but also eliminates the need for separate lookups based on precision levels, thereby enhancing memory access efficiency during attention computations. The memory overhead of the bidirectional page table is minimal: for example, with a batch size of 128 on Llama-2-7B, which has 32 layers and 32 attention heads per layer, the total size of all bidirectional page tables is only 64 MB. By contrast, the KV cache for a single request occupies 2 GB.
5.3 KV Compaction Workflow
@@
We delve into the KV compaction workflow during the prompt and generation phases, illustrating how the three data structures interact to enable differentiated KV cache compression. Notably, KV compaction is executed once per inference step for all requests in the batch and all layers in the LLM. This design ensures sufficient parallelism for efficient GPU execution of parallel KV compaction, while amortizing the associated GPU kernel launch overheads.
Figure 5: Memory management flow in the prompt phase.
@@
Prompt Phase. Figure 5 illustrates the KV compaction workflow during the prompt phase using an example request with eight tokens. In this setting, one high-precision page can store two tokens, and one low-precision page can store four. Since the exact number of high- and low-precision pages required by each head is unknown prior to running the compression algorithm, we conservatively allocate four unified pages per head, assuming that all tokens will be stored at high precision. Each head then independently applies the KV cache compression algorithm to determine which tokens to store at high-precision, low-precision, or to prune. This step constitutes the planning phase, during which each head calculates its specific memory needs based on its unique sparsity pattern. In this example, the first head uses one high-precision page, while the second head uses two; both heads also require one low-precision page. High-precision pages are allocated from the left side of the bidirectional page table, while low-precision pages are allocated from the right side. Following the planning phase, any unused pages are marked for recycling, triggering the coordination phase. Here, we employ a parallel prefix-sum-based approach, as described in the previous section, to facilitate efficient and conflict-free page recycling across heads.
@@
6 Implementation
We implement LeanKV on top of vLLM, comprising 4.5K lines of CUDA/C++ code and 9K lines of Python. We first outline the architecture of LeanKV and then detail our custom GPU attention kernel, designed to efficiently support the differentiated KV cache compression.
6.1 LeanKV Architecture
The architecture of LeanKV is illustrated in Figure 6. At each inference step, the scheduler batches as many requests as possible within the available GPU memory to maximize throughput, sending the selected requests to all workers. Each GPU hosts one worker, responsible for executing a segment of the model. LeanKV utilizes tensor parallelism [54], where each worker handles a partition of each matrix multiplication and a subset of the attention heads. Each worker includes a dedicated memory manager to oversee the KV cache for its assigned attention heads, using the method described in Section 5. Due to dynamic sparsity patterns, memory requirements vary per head for each request, leading to an irregular memory layout. Additionally, each worker incorporates an execution engine for model computation. To support differentiated KV cache compression, the execution engine integrates a KV compressor and a custom GPU attention kernel. After computing key and value vectors, the KV compressor is invoked to compress them following the policy described in Section 4, storing the results in the KV cache. The custom GPU attention kernel then efficiently computes the attention output using the compressed KV cache.
Figure 6: LeanKV architecture.
6.2 Efficient Attention Kernel
We develop a custom GPU attention kernel that efficiently supports differentiated KV cache compression, fully harnessing the reduction in memory access volume and translating it into accelerated performance.
Overall, we assign each CUDA thread block to process a single attention head per sequence. However, mixed-precision quantization introduces potential load imbalance, as high-precision tokens demand more bandwidth, which is the bottleneck of attention, than low-precision ones. Thanks to unified pages that store tokens of the same precision together, we mitigate this issue by having thread warps iterate over high-precision pages first, followed by low-precision pages, ensuring that threads within a warp remain load-balanced. Each page is processed in two phases: dot product between the query and keys to derive attention scores, and weighted sum of values. These two phases follow different computation patterns: reduction over the feature dimension and token dimension, respectively. To ensure coalesced and vectorized memory accesses for both phases, we design tailored data layouts and parallelization strategies, elaborated as follows.